# Knowledge Editing for Large Language Models: A Survey

### v3

@@ 3. Problem Formulation
Editing Target. In this survey, we represent the knowledge encoded or required to be injected into the pre-trained LLMs as a knowledge triple $t=(s,r,o)$, where $s$ is the subject (e.g., president of the USA), $r$ is the relation (e.g., is), and $o$ is the object (e.g., Biden). From the perspective of knowledge triple, the objective of KME for LLMs is to modify the original knowledge triple $t=(s,r,o)$ encoded in the pre-trained weights of the model into the target knowledge triple $t^{*}=(s,r,o^{*})$, where $o^{*}$ is the target object different from $o$. In this manner, we can define an edit as a tuple $e=(t,t^{*})=(s,r,o\rightarrow o^{*})$, which denotes the update of the obsolete old knowledge $t$ into the new knowledge $t^{*}$.
@@
Input and Output Space. Given a pair of subject $s$ and relation $r$, in order to query LLMs to obtain the object $o$, $(s,r)$ needs to be transformed into natural language, which we denoted as $x$. $x$ is also referred to as the prompt in this survey. The LLM output $y$ is also textual and can be converted back to an object $o$ as the query result. In this manner, $(x,y)$ can be considered as the natural language input-output pair associated with the knowledge triple $t=(s,r,o)$. For example, the prompt $x$ transformed from $s$ and $r$ can be “The president of the USA is”, and $y$ is the model output “Joe Biden”. Note that due to the diversity of natural language, multiple $(x,y)$ pairs can be associated with the same knowledge triple $t$. We denote the set of textual inputs that can be associated with subject $s$ and relation $r$ in an edit $e$ as $\mathcal{X}_{e}=I(s,r)$, referred to as in-scope input space. Similarly, we define the set of textual outputs that can be associated with the object $o$ in the same edit $e$ as $\mathcal{Y}_{e}^{*}=O^{*}(s,r,o^{*})$ (i.e., target output space), and the original textual output space as $\mathcal{Y}_{e}=O(s,r,o)$ (i.e., original output space). Given an edit $e$, the aim of KME is to modify the behavior of language models from $\mathcal{Y}_{e}$ to $\mathcal{Y}_{e}^{*}$, regarding the input in $\mathcal{X}_{e}$. To accommodate the scenarios where multiple edits are performed, we can define the union of $\mathcal{X}_{e}$ over a set of edits =$\mathcal{E}=\{e_{1},e_{2},\dotsc\}$ as $\mathcal{X}_{\mathcal{E}}=\bigcup_{e\in\mathcal{E}}\mathcal{X}_{e}$. Similarly, we can define $\mathcal{Y}_{\mathcal{E}}=\bigcup_{e\in\mathcal{E}}\mathcal{Y}_{e}$ and $\mathcal{Y}^{*}_{\mathcal{E}}=\bigcup_{e\in\mathcal{E}}\mathcal{Y}^{*}_{e}$.
@@
[…]
@@ 4.1. Accuracy
Accuracy is the most straightforward measure for the effectiveness of KME techniques, which is defined as the success rate of editing in terms of a specific set of pre-defined input-output pairs =$\mathcal{E}=\{x_{i},y^{*}_{i}\}$ associated with all the edited knowledge. Accuracy can be easily defined to evaluate the performance of KME on classification tasks, e.g., fact checking (Mitchell et al., 2022b; Petroni et al., 2021), where the answers $y$ are categorical. If we define the prompt and ground truth related to an edit $e$ as $x_{e}$ and $y^{*}_{e}$, respectively, the metric of the accuracy of an edited model $f^{*}$ can be formulated as follows:
@@
Since accuracy is defined on a deterministic set of prompts-answer pairs =$\mathcal{E}=\{x_{i},y^{*}_{i}\}$, it provides a fair comparison between KME methods. Nevertheless, it is challenging to evaluate the practicality of KME methods with accuracy, as there is no consensus on how to design the $\mathcal{E}$, especially when the task needs to output a long sequence such as question answering or text generation (Meng et al., 2022; Dong et al., 2022a; Meng et al., 2023).
@@ 4.2. Locality
One crucial property for the KME strategies is locality, which measures the capability of the edited model $f^{*}$ to preserve the pre-trained knowledge in $f$ irrelevant to the edits in $\mathcal{E}$. We note that in most cases of KME, the number of required edits makes for an extremely small fraction of the entire knowledge learned and preserved in the pre-trained LLMs. Consequently, the locality measurement is of great importance in assessing the capability of edited models to preserve unrelated knowledge. Given an edit $e$, an edited model $f^{*}$, and the original pre-trained model $f$, the locality of $f^{*}$ can be defined as the expectation of unmatched agreement between the edited model and unedited model for out-scope inputs, which can be defined as follows:
(4) $\textbf{Loc}(f^{*},f;e)=\mathbb{E}_{x\notin\mathcal{X}_{e}}\mathbbm{1}\{f^{*}(x)\neq f(x)\}.$
@@
(5) $\textbf{Loc}(f^{*},f;\mathcal{E})=\mathbb{E}_{x\notin\mathcal{X}_{\mathcal{E}}}\mathbbm{1}\{f^{*}(x)\neq f(x)\},\ \ \text{where}\ \ \mathcal{X}_{\mathcal{E}}=\bigcup_{e\in\mathcal{E}}\mathcal{X}_{e}.$
Although the above metric measures the overall locality of $f^{*}$ based on all inputs that are not in $\mathcal{X}_{\mathcal{E}}$, it is difficult to compute in realistic scenarios, as the entire input space can be excessively large or even infinite. Therefore, existing methods generally resort to alternative solutions that pre-define the specific range of out-scope inputs to calculate the expected difference in Eqs. (4) and (5). For example, in SERAC (Mitchell et al., 2022b), the authors generate hard out-scope examples from the dataset zsRE (Levy et al., 2017) by selectively sampling from training inputs with high semantic similarity with the edit input, based on embeddings obtained from a pre-trained semantic embedding model. If we denote the out-scope input space related to the input $\mathcal{X}_{e}$ as $\mathcal{O}_{e}$, we can similarly define the feasible out-scope input space for multiple edits as $\mathcal{O}_{\mathcal{E}}=\bigcup_{e\in\mathcal{E}}\mathcal{O}_{e}$. In this manner, we define a specific metric of locality of $f^{*}$ regarding $\mathcal{E}$ as follows:
(6) $\textbf{Loc}(f^{*},f;\mathcal{O}_{e})=\mathbb{E}_{x\in\mathcal{O}_{e}}\mathbbm{1}\{f^{*}(x)\neq f(x)\},$
(7) $\textbf{Loc}(f^{*},f;\mathcal{O}_{\mathcal{E}})=\mathbb{E}_{x\in\mathcal{O}_{\mathcal{E}}}\mathbbm{1}\{f^{*}(x)\neq f(x)\},\ \ \text{where}\ \ \mathcal{O}_{\mathcal{E}}=\bigcup_{e\in\mathcal{E}}\mathcal{O}_{e}.$
@@ 4.3. Generality
Aside from locality, another crucial property to measure the effectiveness of KME is generality, which indicates the capability of the edited model $f^{*}$ to correctly respond to similar prompts. This requires the generalization of the updated knowledge to other in-scope inputs that do not appear in the training set while conveying similar or related meanings. Ensuring the generality of edited models prevents the edited model from overfitting to a particular input. Specifically, in the scenarios of knowledge-based model editing, the inherent diversity of natural language determines that various in-scope inputs $x$ can correspond to a specific knowledge triple $t$ (Wang et al., 2023b). These semantically equivalent inputs can involve differences in aspects such as syntax, morphology, genre, or even language. Existing works mostly pre-define a specific in-scope input space of each edit via different strategies. For example, in the CounterFact dataset proposed in ROME (Meng et al., 2022), the authors utilize prompts that involve distinct yet semantically related subjects as the in-scope input. In general, the generality of an edited model $f^{*}$ is defined as the expectation of exact-match agreement between the output of the edited model and true labels for in-scope inputs, which can be defined on either an edit $e$ or the edit set $\mathcal{E}$ as follows:
@@ 4.4. Retainability
Retainability characterizes the ability of KME techniques to preserve the desired properties of edited models after multiple consecutive edits. In the presence of ever-evolving information, practitioners may need to frequently update a conversational model (i.e., sequential editing). Such a KME setting requires that the model not forget previous edits after each new modification. It is essential to distinguish retainability from scalability, which evaluates the model’s ability to handle a vast number of edits. In contrast, retainability assesses the consistent performance of the model after each individual edit, presenting a more challenging objective to achieve. Recently, T-Patcher (Huang et al., 2023) first explores the sequential setting of KME and observes that many existing approaches significantly fall short in terms of retainability. In SLAG (Hase et al., 2023), the authors also discover a significant drop in editing performance when multiple beliefs are updated continuously. To thoroughly assess the retainability of an edited language model $f^{*}$, we define it as follows:
@@ 4.5. Scalability
The scalability of an editing strategy refers to its capability to incorporate a large number of edits simultaneously. Recently, several works have emerged that can inject multiple new knowledge into specific parameters of pre-trained LLMs. For instance, SERAC (Mitchell et al., 2022b) can perform a maximum of 75 edits. In addition, MEMIT (Meng et al., 2023) is proposed to enable thousands of edits without significant influence on editing accuracy. When there is a need to edit a model with a vast number of edits concurrently, simply employing the current knowledge-based model editing techniques in a sequential manner is proven ineffective in achieving such scalability (Yao et al., 2023). To effectively evaluate the scalability of edited language models, we define the scalability of an edited model as follows:
@@
Figure 3. The detailed categorization of knowledge-based model editing (KME) techniques for LLMs.
Table 2. Comparisons among different categories of methods across the four metrics. The properties that each type of strategy focuses on are highlighted with check marks.
Category Strategy Locality Generality Retainability Scalability References
External Memorization Memory-based (Mitchell et al., 2022b; Murty et al., 2022; Madaan et al., 2022; Li et al., 2023c; Zheng et al., 2023; Zhong et al., 2023)
Extension-based (Dong et al., 2022a; Huang et al., 2023; Hartvigsen et al., 2022)
Global Optimization Constrained Fine-tuning (Zhu et al., 2020; Chen et al., 2020; Lee et al., 2022; Sinitsin et al., 2020; Ni et al., 2023),
Intermediate Fine-tuning (Ha et al., 2016; De Cao et al., 2021; Hase et al., 2023; Mitchell et al., 2022a; Cheng et al., 2023b)
Local Modification Gradient-based (Dai et al., 2022; Meng et al., 2022, 2023; Li et al., 2023b; Yu et al., 2023; Wu et al., 2023)
Prompt-based (Meng et al., 2022; Gupta et al., 2023; Ma et al., 2023)

### v4

@@ 3. Problem Formulation
Editing Target. In this survey, we represent the knowledge required to be injected into LLMs as a knowledge triple $t=(s,r,o)$, where $s$ is the subject (e.g., president of the USA), $r$ is the relation (e.g., is), and $o$ is the object (e.g., Biden). From the perspective of knowledge triple, the objective of KME for LLMs is to modify the original knowledge triple $t=(s,r,o)$ encoded in the pre-trained weights of the model into the target knowledge triple $t^{*}=(s,r,o^{*})$, where $o^{*}$ is the target object different from $o$. In this manner, we can define an edit as a tuple $e=(t,t^{*})=(s,r,o\rightarrow o^{*})$, which denotes the update of the obsolete old knowledge $t$ into the new knowledge $t^{*}$.
@@
Input and Output Space. Given a pair of subject $s$ and relation $r$, in order to query LLMs to obtain the object $o$, $(s,r)$ needs to be transformed into natural language, which we denoted as $x$. $x$ is also referred to as the prompt in this survey. The LLM output $y$ is also textual and can be converted back to an object $o$ as the query result. In this way, $(x,y)$ can be considered as the natural language input-output pair associated with the knowledge triple $t=(s,r,o)$. For example, the prompt $x$ transformed from $s$ and $r$ can be “The president of the USA is”, and $y$ is the model output “Joe Biden”. Note that due to the diversity of natural language, multiple $(x,y)$ pairs can be associated with the same knowledge triple $t$. We denote the set of textual inputs associated with subject $s$ and relation $r$ in an edit $e$ as $\mathcal{X}_{e}=I(s,r)$, referred to as in-scope input space. Similarly, we define the set of textual outputs that can be associated with the object $o$ in the same edit $e$ as $\mathcal{Y}_{e}^{*}=O^{*}(s,r,o^{*})$ (i.e., target output space), and the original textual output space as $\mathcal{Y}_{e}=O(s,r,o)$ (i.e., original output space). Given an edit $e$, the aim of KME is to modify the behavior of language models from $\mathcal{Y}_{e}$ to $\mathcal{Y}_{e}^{*}$, regarding the input in $\mathcal{X}_{e}$. To accommodate the scenarios where multiple edits are performed, we can define the union of $\mathcal{X}_{e}$ over a set of edits =$\mathcal{E}=\{e_{1},e_{2},\dotsc\}$ as $\mathcal{X}_{\mathcal{E}}=\bigcup_{e\in\mathcal{E}}\mathcal{X}_{e}$. Similarly, we can define $\mathcal{Y}_{\mathcal{E}}=\bigcup_{e\in\mathcal{E}}\mathcal{Y}_{e}$ and $\mathcal{Y}^{*}_{\mathcal{E}}=\bigcup_{e\in\mathcal{E}}\mathcal{Y}^{*}_{e}$.
@@
(2) $\min\mathbb{E}_{e\in\mathcal{E}}\mathbb{E}_{x,y^{*}\in\mathcal{X}_{e},\mathcal{Y}^{*}_{e}}\mathcal{L}(f^{*}(x),y^{*}),\text{where}\ \ f^{*}=M(f;\mathcal{E}),$
@@ 4.1. Accuracy
Accuracy is a straightforward metric for evaluating the effectiveness of KME techniques (Li et al., 2023b; Zheng et al., 2023; Zhong et al., 2023; Dong et al., 2022a; Ni et al., 2023; Mitchell et al., 2022a; Cheng et al., 2024), defined as the success rate of editing in terms of a specific set of pre-defined input-output pairs $(x_{e},y^{*}_{e})$ associated with all the edited knowledge. Accuracy can be easily defined to evaluate the performance of KME on classification tasks, e.g., fact checking (Mitchell et al., 2022b; Petroni et al., 2021), where the answers $y$ are categorical. Defining the prompt and ground truth related to an edit $e$ as $x_{e}$ and $y^{*}_{e}$, respectively, the metric of the accuracy of an edited model $f^{*}$ is formulated as follows:
@@
Since accuracy is defined on a deterministic set of prompt-answer pairs, it provides a fair comparison between KME methods (Dai et al., 2022; Meng et al., 2022, 2023). Nevertheless, it is non-trivial to evaluate the practicality of KME methods with accuracy, as there is no consensus on how to design the $\mathcal{E}$, especially when the task needs to output a long sequence such as question answering or text generation (Meng et al., 2022; Dong et al., 2022a; Meng et al., 2023).
@@ 4.2. Locality
One crucial metric for the KME strategies is locality (De Cao et al., 2021; Mitchell et al., 2022a; Cheng et al., 2024; Li et al., 2024c), which reflects the capability of the edited model $f^{*}$ to preserve the pre-trained knowledge in $f$ irrelevant to the edits in $\mathcal{E}$. Note that in most KME applications, the number of required edits makes for an extremely small fraction of the entire knowledge learned and preserved in the pre-trained LLMs (Yao et al., 2023; Zhang et al., 2024). Consequently, the locality measurement is of great importance in assessing the capability of edited models to preserve unrelated knowledge (Murty et al., 2022; Madaan et al., 2022; Gupta et al., 2023). Given an edit $e$, the edited model $f^{*}$, and the original pre-trained model $f$, the locality of $f^{*}$ can be defined as the expectation of matched agreement between the edited model and unedited model for out-scope inputs, which can be defined as follows:
(4) $\textbf{Loc}(f^{*},f;e)=\mathbb{E}_{x\notin\mathcal{X}_{e}}\mathbbm{1}\{f^{*}(x)=f(x)\}.$
@@
(5) $\textbf{Loc}(f^{*},f;\mathcal{E})=\mathbb{E}_{x\notin\mathcal{X}_{\mathcal{E}}}\mathbbm{1}\{f^{*}(x)=f(x)\},\ \ \text{where}\ \ \mathcal{X}_{\mathcal{E}}=\bigcup_{e\in\mathcal{E}}\mathcal{X}_{e}.$
Although the above metric measures the overall locality of $f^{*}$ based on all inputs that are not in $\mathcal{X}_{\mathcal{E}}$, it is difficult to compute in realistic scenarios, as the entire input space can be excessively large or even infinite (Yao et al., 2023). Therefore, existing methods generally resort to alternative solutions that pre-define the specific range of out-scope inputs to calculate the locality metric (De Cao et al., 2021; Dai et al., 2022; Meng et al., 2022; Chen et al., 2024; Li et al., 2024b). For example, in SERAC (Mitchell et al., 2022b), the authors generate hard out-scope examples from the dataset zsRE (Levy et al., 2017) by selectively sampling from training inputs with high semantic similarity with the edit input, based on embeddings obtained from a pre-trained semantic embedding model. Denoting the out-scope input space related to the input $\mathcal{X}_{e}$ as $\mathcal{O}_{e}$, we can similarly define the feasible out-scope input space for multiple edits as $\mathcal{O}_{\mathcal{E}}=\bigcup_{e\in\mathcal{E}}\mathcal{O}_{e}$. In this manner, we define a specific metric of locality of $f^{*}$ regarding $\mathcal{E}$ as follows:
(6) $\textbf{Loc}(f^{*},f;\mathcal{O}_{e})=\mathbb{E}_{x\in\mathcal{O}_{e}}\mathbbm{1}\{f^{*}(x)=f(x)\},$
(7) $\textbf{Loc}(f^{*},f;\mathcal{O}_{\mathcal{E}})=\mathbb{E}_{x\in\mathcal{O}_{\mathcal{E}}}\mathbbm{1}\{f^{*}(x)=f(x)\},\ \ \text{where}\ \ \mathcal{O}_{\mathcal{E}}=\bigcup_{e\in\mathcal{E}}\mathcal{O}_{e}.$
@@ 4.3. Generality
Aside from locality, another crucial metric is generality, which indicates the capability of the edited model $f^{*}$ to correctly respond to semantically similar prompts (Zhu et al., 2020; Chen et al., 2020; Sharma et al., 2024; Ni et al., 2023; Mitchell et al., 2022a). This requires the generalization of the updated knowledge to other in-scope inputs that do not appear in the training set while conveying similar or related meanings (Gupta et al., 2024; Wei et al., 2024). As such, ensuring the generality of edited models prevents the edited model from overfitting to a particular input (Zhang et al., 2024). Specifically, in the scenarios of knowledge-based model editing, the inherent diversity of natural language determines that various in-scope inputs $x$ can correspond to a specific knowledge triple $t$ (Wang et al., 2023c). These semantically equivalent inputs can involve differences in aspects such as syntax, morphology, genre, or even language. Existing works mostly pre-define a specific in-scope input space of each edit via different strategies (Yoon et al., 2024; Hu et al., 2024; Wu et al., 2024; Song et al., 2024; Li et al., 2024d). For example, in the CounterFact dataset proposed in ROME (Meng et al., 2022), the authors utilize prompts that involve distinct yet semantically related subjects as the in-scope input. In general, the generality of an edited model $f^{*}$ is defined as the expectation of exact-match agreement between the output of the edited model and true labels for in-scope inputs, which can be defined on either an edit $e$ or the edit set $\mathcal{E}$ as:
@@
4.4. Portability
In addition to generality, another vital metric is portability, which measures the effectiveness of the edited model $f^{*}$ in transferring a conducted edit to other logically related edits that can be interpreted via reasoning (Zhang et al., 2024). For example, if an edit is conducted towards the President of the USA, the edit regarding the query “Which political party does the current President of the USA belong to?” should also be achieved. This ensures that the edited model is not limited to responding to specific input formats. In concrete, the transfer of knowledge is crucial for robust generalization of the edited model. In practice, portability can be assessed with logically related edits obtained in different ways (Yao et al., 2023; Cohen et al., 2024). Denoting an edit as $e=(s,r,o\rightarrow o^{*})$, hereby we introduce two common types of logically related edits $\tilde{e}$. (1) Reversed Relation: $\tilde{e}=(o\rightarrow o^{*},\tilde{r},s)$, where $\tilde{r}$ is the reversed relation of $r$, and (2) Neighboring Relation: $\tilde{e}=(s,r\oplus r_{\epsilon},\epsilon\rightarrow\epsilon^{*})$, where both $(o,r_{\epsilon},\epsilon)$ and $(o^{*},r_{\epsilon},\epsilon^{*})$ exist in the pre-trained knowledge, and $r\oplus r_{\epsilon}$ is a combined relation from $r$ and $r_{\epsilon}$. In this manner, we define portability as the edited model performance on one or multiple logically related edits as follows:
(10) $\textbf{Por}(f^{*};\tilde{e})=\mathbb{E}_{x\in\mathcal{X}_{\tilde{e}}}\mathbbm{1}\{f^{*}(x)\in\mathcal{Y}_{\tilde{e}}^{*}\},$
[…]
@@ 4.5. Retainability
Retainability characterizes the ability of KME techniques to preserve the desired properties of edited models after multiple consecutive edits (Jiang et al., 2024; Yu et al., 2024; Gu et al., 2023). In the presence of ever-evolving information, practitioners may need to frequently update a conversational model (i.e., sequential editing). Such a KME setting requires that the model does not forget previous edits after each new modification (Li et al., 2024a). It is essential to distinguish retainability from scalability, which evaluates the model’s ability to handle a vast number of edits (Chen et al., 2024). In contrast, retainability assesses the consistent performance of the model after each individual edit, presenting a more challenging objective to achieve. Recently, T-Patcher (Huang et al., 2023) first explores the sequential setting of KME and observes that many existing approaches significantly fall short in terms of retainability. In SLAG (Hase et al., 2023), the authors also discover a significant drop in editing performance when multiple beliefs are updated continuously. To assess the retainability of an edited language model $f^{*}$, we define it as follows:
@@ 4.6. Scalability
The scalability of an editing strategy refers to its capability to incorporate a large number of edits simultaneously (Chen et al., 2024). Recently, several works have emerged that can inject multiple new knowledge into specific parameters of pre-trained LLMs (Yoon et al., 2024; Zhang et al., 2024). For instance, SERAC (Mitchell et al., 2022b) can perform a maximum of 75 edits. In addition, MEMIT (Meng et al., 2023) is proposed to enable thousands of edits without significant influence on editing accuracy. When there is a need to edit a model with a vast number of edits concurrently, simply employing the current knowledge-based model editing techniques in a sequential manner is proven ineffective in achieving such scalability (Yao et al., 2023). To effectively evaluate the scalability of edited language models, we define the scalability of an edited model as follows:
@@
[…]
{forest}
for tree= edge=-, draw=connect-line, line width=1pt, grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=middle, font=, rectangle, line width=1.2pt, draw=connect-line, rounded corners,align=left, minimum width=2em, s sep=4pt, inner xsep=3pt, inner ysep=0pt, , where level=1text width=4.5em, where level=2text width=5em,font=, where level=3font=, where level=4straight edge, font=, where level=5font=, [KME , yellow-root ,edge=external, [External Memorization, external-middle, edge=external, text width=7.5em, fill=external!30 [Memory-based, external-middle, text width=8.2em, edge=external, fill=external!30 [ MeLLo (Zhong et al., 2023) , MemPrompt (Madaan et al., 2022), IKE (Zheng et al., 2023),
Language Patch (Murty et al., 2022), SERAC (Mitchell et al., 2022b), KAFT (Li et al., 2023b) , external-leaf, text width=15.8em, edge=external ] ] [Extension-based, external-middle, text width=8.2em, edge=external, fill=external!30 [CALINET (Dong et al., 2022a), T-Patcher (Huang et al., 2023), GRACE (Hartvigsen et al., 2023),
COMEBA-HK (Li et al., 2024a), SWEA (Li et al., 2024b) , external-leaf, text width=15.8em, edge=external] ] ] [Global Optimization, global-middle, edge=global, text width=7.5em, fill=global!30 [Constrained Fine-tuning, global-middle, text width=8.2em, edge=global, fill=global!30 [RecAdam(Chen et al., 2020), Editable Training (Sinitsin et al., 2020),
PPA (Lee et al., 2022), Modifying-Memory(Zhu et al., 2020),
F-Learning (Ni et al., 2023), MELO (Yu et al., 2024), RECT (Gu et al., 2024) ,global-leaf, text width=15.8em, edge=global,] ] [Intermediate Fine-tuning, global-middle, text width=8.2em, edge=global, fill=global!30 [ KGEditor (Cheng et al., 2024), KE (De Cao et al., 2021), SLAG (Hase et al., 2023), MEND(Mitchell et al., 2022a) , global-leaf, text width=15.8em, edge=global] ] ] [Local Modification, local-middle, edge=local, text width=7.5em, fill=local!30 [Groundtruth-based, local-middle, text width=8.2em, edge=local,fill=local!30 [KD(Dai et al., 2022), ROME (Meng et al., 2022), DEPN (Wu et al., 2023), PMET (Li et al., 2024c),
MEMIT (Meng et al., 2023), EMMET (Gupta et al., 2024), DINM (Wang et al., 2024d) , local-leaf, text width=15.8em, edge=local ] ] [Prompt-based, local-middle, text width=8.2em, edge=local,fill=local!30 [MEMITCSK (Gupta et al., 2023), BIRD (Ma et al., 2023) , local-leaf, text width=15.8em, edge=local] ] ] ]
Figure 3. The categorization of KME techniques for LLMs and the corresponding works.